OpenStudy (anonymous):

The defining equation for the function g is g(x) = x^2 - 2x + 3. Find g(3+h)

OpenStudy (anonymous):

9+h 3h^2 - 6h + 6 h^2 + 4h + 6 2x^2 - 2x + h + 52

OpenStudy (gorv):

g(3+h)=(3+h)^2-2*(3+h)+3 =9+h^2+2*3*h-6-2h+3 =h^2+6h-2h+9+3-6 =h^2+4h+6

OpenStudy (gorv):

just replace x by 3+h